# Rotary Position Embedding for Vision Transformer

### v1

@@ 4.2 Object detection
Axial freqs Mixed freqs Mixed + APE
ViT-B 49.4 50.8(+1.4) 51.2(+1.8) 51.1(+1.7)
ViT-L 51.1 52.2(+1.1) 52.9(+1.8) 52.8(+1.7)
Table 2: MS-COCO object detection with DINO Swin. MS-COCO [16] detection performance (box AP) is reported for SwinTransformer with RoPE. DINO [39] is trained with DINO Swin 12 epochs setting [24]. SwinTransformer with RoPE or RPE is pretrained on ImageNet-1k with SwinTransformer 300epochs recipe.
@@
Axial freqs Mixed freqs Mixed + RPB
Swin-T 51.3 51.6(+0.3) 51.8(+0.5) 51.6(+0.3)
Swin-S 53.0 53.1(+0.1) 53.3(+0.3) 53.6(+0.6)
Swin-B 54.2 54.4(+0.2) 54.5(+0.3) 54.5(+0.3)
We verify 2D RoPE in an object detection on MS-COCO dataset [16]. DINO [39] detector is trained using ViT and Swin Transformer as backbone network. We use ImageNet-1k supervised learning from §4.1 for backbone pre-training, and RoPE is only applied to the backbone network. The Detrex [24] codebase is used for detection training. For ViT, we use the DINO-ViTDet 12 epochs training setting, and the DINO-Swin 12 epochs setting is selected for Swin-based DINO.
Table 1 shows the DINO-ViTDet results in bounding box AP. We report three variants of RoPEs: RoPE-Axial, RoPE-Mixed, and RoPE-Mixed + APE; all demonstrate remarkable performance improvements. DINO ViTDet achieves AP improvement of more than +1.0pp by changing positional embedding to RoPE. Among RoPE variants, RoPE-Mixed shows the best improvement at +1.8pp. AP in ViT-B and ViT-L. DINO-ViTDet uses ViT backbone with window-block attention, but still, a few layers remain as global attention. We believe that RoPE is highly effective due to the extrapolation in global attention, where it outperforms other position embeddings.
The performance of DINO-Swin is reported in Table 2. Like DINO-ViTDet, three variants, RoPE-Axial, RoPE-Mixed, and RoPE-Mixed + RPB, are reported. RoPE outperforms RPB for all variants. RoPE-Mixed performs better than RoPE-Axial, and +RPB is only beneficial in Swin-S. Performance improvement is smaller than DINO-ViTDet since DINO-Swin maintains a window size of the pre-trained backbone, i.e., DINO-Swin has no extrapolation. However, RoPE achieves meaningful gains and it has room for improvement by increasing the Swin Transformer’s window size for the detection backbone.
@@ 4.3 Semantic segmentation
Backbone Multi-scale APE RoPE
Axial freqs Mixed freqs Mixed + APE
ViT-B - 47.7 49.0(+1.3) 49.6(+1.9) 50.0(+2.3)
48.4 49.9(+1.5) 50.7(+2.3) 50.9(+2.5)
ViT-L - 50.8 51.8(+1.0) 51.5(+0.7) 52.0(+1.2)
51.6 52.6(+1.0) 52.3(+0.7) 52.6(+1.0)
Table 4: ADE20k semantic segmentation with Swin Mask2Former [2]. Mask2Former model for semantic segmentation is trained using Swin Transformers. The table shows segmentation performance in mIoU metric.
@@
Axial freqs Mixed freqs Mixed + RPB
Swin-S 50.2 50.4(+0.2) 51.1(+0.9) 50.9(+0.7)
Swin-B 51.5 52.0(+0.5) 52.0(+0.5) 51.4(-0.1)
We train 2D RoPE ViT and Swin Transformer for semantic segmentation tasks on ADE20k [40, 41] dataset. For ViT, we use UperNet [37] with ViT backbone training recipe [21]. For Swin, Mask2Former [2] for semantic segmentation is used with the Swin backbone. ImageNet-1k supervised weights from §4.1 are used for backbone weights, the same as the detection setting. Also, RoPE is only applied to the backbone network. ViT-UperNet and Swin-Mask2Former are trained for 160k iterations.
Table 3 shows ViT-UperNet performance with APE and RoPE. RoPE-based models achieve impressive performance improvement in all cases. It is noteworthy that RoPE-Mixed + APE achieves +2.3 and +2.5 mIoU improvement with only position embedding changes. The improvement might originate from the extrapolation performance of RoPE since the ViT-UperNet setting uses $512\times 512$ images for backbone inputs. Among the three variants of 2D RoPE, RoPE-mixed + APE shows the best performance in all cases, which is different from detection results. As shown in Fig. 5, RoPE-mixed + APE has an advantage at interpolation while degrading performance at extrapolation. These results imply that the use of APE in a RoPE-based network should change depending on the target task. Swin-Mask2Former performances are shown in Table 4. RoPE also improves the performance of Swin-based segmentation. RoPE-Mixed shows the best performance among the three 2D RoPE variants. RoPE-Mixed + RPB has failed to improve RoPE Swin’s performance.
@@ 4.4 Comparison with multi-resolution methods
We compare 2D RoPE variants with recent ViT architecture designed for multi-resolution inference, namely ResFormer [30]. ResFormer uses depth-wise convolutions as the position embedding method. It uses sinusoidal APE in Eq. 2 and depth-wise convolution after the patch-embed layer as Global Position Embedding (GPE). Also, another depth-wise convolution is used similar to skip-connection for every self-attention layer to add position embed as Local Position Embed (LPE). Using GPE and LPE, ResFormer is proposed as an improved ViT for multi-resolution inference. ResFormer is trained with multi-resolution training utilizing self-distillation loss. Since self-distillation with multi-resolution training is not a common recipe in ViT, we use ResFormer-S trained with fixed resolution $224\times 224$ and compare it with RoPE-Mixed ViT-S in §4.1. Table 5 shows a multi-resolution comparison of RoPE-Mixed with ResFormer-S-224. RoPE-Mixed outperforms ResFormer with a meaningful margin for extrapolation ranges (res $>224$), but RoPE-Mixed shows performance lower than ResFormer for significant interpolation ranges (res $\leq 160$). To achieve comparable interpolation, RoPE-Mixed needs additional APE. Overall, the results show that RoPE-Mixed+APE outperforms ResFormer-S in multi-resolution inference.
@@ Appendix 0.A Experiments (cont’d)
We demonstrated the performance of 2D RoPE with performance graphs through various input resolutions in the main paper. This appendix provides the entire performance numbers for the multi-resolution experiments. We measured the performance of ViTs and Swin Transformers with default position embedding (APE or RPB), 2D RoPE variants (RoPE-Axial and RoPE-Mixed), and 2D RoPE variants with default position embedding (RoPE-Axial and RoPE-Mixed + APE or RPB). Note that figures in the paper do not include small resolutions such as $96\times 96$ to improve the visualization.
0.A.1 Multi-resolution classification – ViT
Table 0.A.1, 0.A.2, and 0.A.3 report the total numbers of multi-resolution classification, which is illustrated in Figure 3. 2D RoPE variants outperform APE in the smallest resolution $96\times 96$ with significant gap.
Table 0.A.1: Multi-resolution performance of ViT-S. Table reports the performance of 2D RoPE variants corresponding to the first graph in Figure 3.
@@ 0.A.1 Multi-resolution classification – ViT
Table 0.A.2: Multi-resolution performance of ViT-B. Table reports the performance of 2D RoPE variants corresponding to the second graph in Figure 3.
@@
Table 0.A.3: Multi-resolution performance of ViT-L. Table reports the performance of 2D RoPE variants corresponding to the third graph in Figure 3.
@@
0.A.2 Multi-resolution classification – Swin Transformer
Table 0.A.4, 0.A.5, and 0.A.6 show the total numbers of multi-resolution classification of Swin Transformer with 2D RoPE variants corresponding to Figure 4 of paper. Similar to ViT cases, 2D RoPE variants significantly outperform RPB in small resolutions: $96\times 96$ and $128\times 128$.
Table 0.A.4: Multi-resolution performance of Swin-T. Table reports Swin-T performance with 2D RoPE variants corresponding to the first graph in Figure 4.
@@ 0.A.2 Multi-resolution classification – Swin Transformer
Table 0.A.5: Multi-resolution performance of Swin-S. Table reports Swin-S performance with 2D RoPE variants corresponding to the second graph in Figure 4.
@@
Table 0.A.6: Multi-resolution performance of Swin-B. Table reports Swin-B performance with 2D RoPE variants corresponding to the third graph in Figure 4.
@@
0.A.3 2D RoPE with APE or RPB
In Figure 5 of the paper, we report the performance improvement of RoPE-Mixed compared to base position embeddings: APE or RPB. We provide numbers for Figure 5 in Table 0.A.7 and 0.A.8. Note that each number means performance improvement (%p.) compared to base position embeddings (APE or RPB).
Table 0.A.7: Performance improvement compared to APE in ViT-B. Table shows the improvement over APE, which is shown in the left graph of Figure 5.
@@
Table 0.A.8: Performance improvement compared to RPB in Swin-B. Table shows the improvement over RPB, which is shown in the right graph of Figure 5.

### v2

@@ 4.2 Object detection
Axial Mixed Axial+APE Mixed+APE
ViT-B 49.4 50.8(+1.4) 51.2(+1.8) 50.7(+1.3) 51.1(+1.7)
ViT-L 51.1 52.2(+1.1) 52.9(+1.8) 52.5(+1.4) 52.8(+1.7)
Table 2: MS-COCO object detection with DINO-Swin. MS-COCO [16] detection performance (box AP) is reported for Swin Transformer with RoPE. DINO [39] is trained with DINO Swin 12 epochs setting [24]. Swin Transformers with RoPE or RPE are pre-trained on ImageNet-1k with Swin Transformer 300epochs recipe.
@@
Axial Mixed Axial+RPB Mixed+RPB
Swin-T 51.3 51.6(+0.3) 51.8(+0.5) 51.7(+0.4) 51.6(+0.3)
Swin-S 53.0 53.1(+0.1) 53.3(+0.3) 53.5(+0.5) 53.6(+0.6)
Swin-B 54.2 54.4(+0.2) 54.5(+0.3) 54.7(+0.5) 54.5(+0.3)
We verify 2D RoPE in object detection on MS-COCO [16]. DINO [39] detector is trained using ViT and Swin as backbone network. We use ImageNet-1k weights from §4.1 for pre-trained weights, and RoPE is only applied to the backbone. We use Detrex [24] codebase for detection training. DINO-ViTDet 12 epochs setting and DINO-Swin 12 epochs setting are used for DINO training.
Table 1 shows the DINO-ViTDet results in bounding box AP. We report four variants of RoPEs: Axial, Mixed, Axial+APE, and Mixed+APE; all demonstrate remarkable performance improvements. DINO-ViTDet achieves AP improvement of more than +1.0pp by changing positional embedding to RoPE. Among RoPE variants, RoPE-Mixed shows the best improvement at +1.8pp. AP in ViT-B and ViT-L. DINO-ViTDet uses ViT backbone with window-block attention, but still, a few layers remain as global attention. We believe that RoPE is highly effective due to the extrapolation in global attention.
The performance of DINO-Swin is reported in Table 2. Like DINO-ViTDet, four variants are reported: Axial, Mixed, Axial+RPB, and Mixed+RPB. RoPE outperforms RPB for all variants. RoPE-Mixed performs better than RoPE-Axial. +RPB is beneficial for Axial but has limited effect on Mixed. Performance improvement is smaller than DINO-ViTDet since DINO-Swin maintains a window size of the pre-trained backbone, i.e., DINO-Swin has no extrapolation. However, RoPE achieves meaningful gains and has room for improvement by increasing the Swin Transformer’s window size for the detection backbone.
@@ 4.3 Semantic segmentation
Multi- scale APE RoPE
Axial Mixed Axial+APE Mixed+APE
ViT-B - 47.7 49.0(+1.3) 49.6(+1.9) 49.5(+1.8) 50.0(+2.3)
48.4 49.9(+1.5) 50.7(+2.3) 50.5(+2.1) 50.9(+2.5)
ViT-L - 50.8 51.8(+1.0) 51.5(+0.7) 51.6(+0.8) 52.0(+1.2)
51.6 52.6(+1.0) 52.3(+0.7) 52.4(+0.8) 52.6(+1.0)
Table 4: ADE20k semantic segmentation with Swin-Mask2Former [2]. Mask2Former model for semantic segmentation is trained using Swin Transformer. The table shows segmentation performance in mIoU metric.
@@
Axial Mixed Axial+RPB Mixed+RPB
Swin-S 50.2 50.4(+0.2) 51.1(+0.9) 51.2(+1.0) 50.9(+0.7)
Swin-B 51.5 52.0(+0.5) 52.0(+0.5) 50.0(-1.5) 51.4(-0.1)
We train 2D RoPE ViT and Swin for semantic segmentation on ADE20k [40, 41]. For ViT, we use UperNet [37] with ViT training recipe [21]. For Swin, Mask2Former [2] for segmentation is used with the Swin. ImageNet-1k pre-trained weights from §4.1 are used for pre-trained weights. Also, RoPE is only applied to the backbone. The networks are trained for 160k iterations.
Table 3 shows ViT-UperNet performances. RoPE-based models achieve impressive performance improvement in all cases. It is noteworthy that Mixed+APE achieves +2.3 and +2.5 mIoU improvement with only position embedding changes. The improvement might originate from the extrapolation performance of RoPE since the ViT-UperNet setting uses $512\times 512$ images for inputs. Among the three variants of RoPE, Mixed+APE shows the best performance in all cases, which is different from detection results. As shown in Fig. 6, Mixed+APE has an advantage at interpolation while degrading performance at extrapolation. These results suggest that the use of APE in a RoPE-based ViT should be adjusted based on the target task. Swin-Mask2Former performances are shown in Table 4. RoPE also improves the performance of Swin-based segmentation. RoPE-Mixed shows impressive performance, while +RPB is only beneficial in limited cases.
@@ 4.4 Comparison with multi-resolution methods
We compare 2D RoPE variants with recent ViT architecture designed for multi-resolution inference, namely ResFormer [30]. ResFormer uses depth-wise convolutions as the position embedding. It uses sinusoidal APE in Eq. 2 and depth-wise convolution after the patch-embed layer as Global Position Embedding (GPE). Also, another depth-wise convolution is used similar to skip-connection for every self-attention layer to add position embed as Local Position Embed (LPE). Using GPE and LPE, ResFormer is proposed as an improved ViT for multi-resolution inference. ResFormer is trained with multi-resolution training utilizing self-distillation loss. Since self-distillation with multi-resolution training is not a common recipe in ViT, we use ResFormer-S trained with fixed resolution $224\times 224$ and compare it with RoPE-Mixed ViT-S in §4.1. Table 5 shows a multi-resolution comparison of RoPE-Mixed with ResFormer-S-224. RoPE-Mixed outperforms ResFormer with a meaningful margin for extrapolation ranges (res $>224$), but RoPE-Mixed shows performance lower than ResFormer for significant interpolation ranges (res $\leq 160$). To achieve comparable interpolation, RoPE-Mixed needs additional APE. Overall, the results show that RoPE-Mixed+APE outperforms ResFormer-S in multi-resolution inference.
@@ Appendix 0.A Experiments (cont’d)
We demonstrated the performance of 2D RoPE with performance graphs through various input resolutions in the main paper. This appendix provides additional ablation studies and the entire performance numbers for the multi-resolution experiments. We measured the performance of ViTs and Swin Transformers with default position embedding (APE or RPB), 2D RoPE variants (RoPE-Axial and RoPE-Mixed), and 2D RoPE variants with default position embedding (RoPE-Axial and RoPE-Mixed + APE or RPB). Note that figures in the paper do not include small resolutions such as $96\times 96$ to improve the visualization.
0.A.1 Impacts of learnable frequencies
Studies on applying RoPE to ViT [7, 18, 19] have not considered the learnable frequencies. However, a comparison with RoPE-Axial + learnable frequencies can be an interesting ablation study by revealing the contribution of learnable frequencies on RoPE-Mixed. Table 0.A.1 and Table 0.A.2 show learnable RoPE-Axial performance compared to fixed RoPE-Axial and RoPE-Mixed. The learnable frequencies improve RoPE-Axial on high-resolution (384, 512) but are ineffective on other resolutions. These results imply that the effects of RoPE-Mixed originate from frequency mixing rather than frequency learning, as we claimed in the paper.
Table 0.A.1: RoPE-Axial with learnable frequencies for ViT-S.
Resolution 144 192 224 256 320 384 512
Axial 73.6 79.2 80.9 81.7 81.5 80.0 76.1
Axial+learn 73.5 79.1 80.7 81.5 81.8 81.3 77.8
(-0.1) (-0.1) (-0.2) (-0.2) (+0.3) (+1.3) (+1.7)
Mixed 74.2 79.6 80.9 81.8 82.2 81.8 79.1
(+0.6) (+0.4) (0.0) (+0.1) (+0.7) (+1.8) (+3.0)
Table 0.A.2: RoPE-Axial with learnable frequencies for ViT-B.
Resolution 144 192 224 256 320 384 512
Axial 78.9 82.8 83.6 84.2 84.3 83.9 82.0
Axial+learn 78.9 82.7 83.6 84.2 84.4 83.9 82.3
(0.0) (-0.1) (0.0) (0.0) (+0.1) 0.0 (+0.3)
Mixed 79.4 82.8 83.8 84.3 84.7 84.4 82.9
(+0.5) (0.0) (+0.2) (+0.1) (+0.4) (+0.5) (+0.9)
0.A.2 Multi-resolution classification – ViT
Table 0.A.3, 0.A.4, and 0.A.5 report the total numbers of multi-resolution classification, which is illustrated in Figure 4. 2D RoPE variants outperform APE in the smallest resolution $96\times 96$ with significant gap.
Table 0.A.3: Multi-resolution performance of ViT-S. Table reports the performance of 2D RoPE variants corresponding to the first graph in Figure 4.
@@ 0.A.2 Multi-resolution classification – ViT
Table 0.A.4: Multi-resolution performance of ViT-B. Table reports the performance of 2D RoPE variants corresponding to the second graph in Figure 4.
@@
Table 0.A.5: Multi-resolution performance of ViT-L. Table reports the performance of 2D RoPE variants corresponding to the third graph in Figure 4.
@@
0.A.3 Multi-resolution classification – Swin Transformer
Table 0.A.6, 0.A.7, and 0.A.8 show the total numbers of multi-resolution classification of Swin Transformer with 2D RoPE variants corresponding to Figure 5 of paper. Similar to ViT cases, 2D RoPE variants significantly outperform RPB in small resolutions: $96\times 96$ and $128\times 128$.
Table 0.A.6: Multi-resolution performance of Swin-T. Table reports Swin-T performance with 2D RoPE variants corresponding to the first graph in Figure 5.
@@ 0.A.3 Multi-resolution classification – Swin Transformer
Table 0.A.7: Multi-resolution performance of Swin-S. Table reports Swin-S performance with 2D RoPE variants corresponding to the second graph in Figure 5.
@@
Table 0.A.8: Multi-resolution performance of Swin-B. Table reports Swin-B performance with 2D RoPE variants corresponding to the third graph in Figure 5.
@@
0.A.4 2D RoPE with APE or RPB
In Figure 6 of the paper, we report the performance improvement of RoPE-Mixed compared to base position embeddings: APE or RPB. We provide numbers for Figure 6 in Table 0.A.9 and 0.A.10. Note that each number means performance improvement (%p.) compared to base position embeddings (APE or RPB).
Table 0.A.9: Performance improvement compared to APE in ViT-B. Table shows the improvement over APE, which is shown in the left graph of Figure 6.
@@
Table 0.A.10: Performance improvement compared to RPB in Swin-B. Table shows the improvement over RPB, which is shown in the right graph of Figure 6.